### v1

# Explicit Lipschitz Value Estimation Enhances Policy Robustness Against Perturbation

Abstract
In robotic control tasks, policies trained by reinforcement learning (RL) in simulation often experience a performance drop when deployed on physical hardware, due to modeling error, measurement error, and unpredictable perturbations in the real world. Robust RL methods account for this issue by approximating a worst-case value function during training, but they can be sensitive to approximation errors in the value function and its gradient before training is complete. In this paper, we hypothesize that Lipschitz regularization can help condition the approximated value function gradients, leading to improved robustness after training. We test this hypothesis by combining Lipschitz regularization with an application of Fast Gradient Sign Method to reduce approximation errors when evaluating the value function under adversarial perturbations. Our empirical results demonstrate the benefits of this approach over prior work on a number of continuous control benchmarks.
1 Introduction
Recently, reinforcement learning (RL) has demonstrated impressive capabilities in robotics control tasks. In practice, RL policies are often trained in physics-based simulators before being transferred to the real world. However, the discrepancy between the simulation and reality, known as the reality gap, can hamper performance when simulated policies are deployed in the real world. Common discrepancies include system identification errors (incorrect mass, friction coefficients, etc. in simulation), idealized physical modeling (e.g., perfectly rigid bodies), and state estimation errors (e.g. sensor noise) in the real world. These discrepancies can be viewed as small differences in the transition dynamics of the simulator vs. those of the real world (Pinto et al., 2017).
Two approaches to bridge the reality gap are to improve the fidelity of the simulators, and to improve the robustness of RL-trained policies to perturbations in the transition dynamics. While some works focus on the first approach, building simulators that more faithfully replicate real-world physics (Tan et al., 2018), a widely used technique in the second approach is domain randomization (DR) (Tobin et al., 2017; Peng et al., 2018; Siekmann et al., 2021). DR works with existing simulators by randomizing the physical parameters (mass, friction, etc.), which effectively perturbs the transition dynamics and exposes the policies to diverse environments during training.
Robust reinforcement learning (Robust RL) uses adversarial perturbations rather than randomly sampled perturbations during training (Iyengar, 2005). This promotes acceptable performance even in a worst case scenario where reality differs from the simulated environment in a pathological way. The trade-off is that best- or average-case performance may be slightly worse compared to policies trained with domain randomization, but this is a small price to pay in high-risk deployment scenarios where dangerous operating modes must be avoided at all costs. Robust RL requires a sub-routine that can approximate the worst-case perturbation at a given state. These approximations are often made using the current value function estimate (Kuang et al., 2022), in which case the training process is sensitive to errors in the value function estimates before learning has converged.
Lipschitz regularization is an effective strategy for reducing sensitivity of neural networks to adversarial examples. It has been used in both computer vision (Miyato et al., 2018) and in RL, to smooth the policy network actions (Song et al., 2023) as well as critic network outputs (Kobayashi, 2022). However, to our knowledge, existing work has not yet studied the efficacy of Lipschitz regularization for reducing sensitivity to value estimates during robust RL training.
In this paper, we show that Lipschitz regularization during robust RL training can improve the robustness of the trained policies in a number of continuous control benchmarks. When generating the adversarial perturbations during training, we incorporate a novel use of Fast Gradient Sign Method (FGSM) (Kurakin et al., 2018) to further reduce errors in value function estimates. Unlike past work that uses FGSM after training (Lütjens et al., 2020) or implicitly as part of a Taylor approximation (Kuang et al., 2022), we re-evaluate the value function on the explicitly computed adversarial perturbation to remove another source of approximation error.
2 Related Work
2.1 Robust Reinforcement Learning
Robust RL (Iyengar, 2005) has drawn increasing attention in recent years, with the advent of adversarial examples and emerging AI safety concerns. Adversarial examples can be incorporated after (Lütjens et al., 2020) or during (Roy et al., 2017; Pinto et al., 2017; Wang & Zou, 2022; Wang et al., 2023) RL training to improve robustness against the worst-case transition dynamics within some compact set of possible environments. Robust RL is relevant to the reality gap as well as other unforeseen circumstances that may arise after deployment. These include unintentional accidents in safety-critical application domains such as power grids and autonomous vehicles (Ilahi et al., 2021), and intentional deceptions launched by malicious actors (Lim et al., 2013; Tessler et al., 2019; Gleave et al., 2019; Zhang et al., 2021). In the latter case, the robust RL optimization is treated as a zero-sum game between the agent and an adversarial opponent (Tessler et al., 2019; Gleave et al., 2019). Usually, the size of adversarial perturbations is constrained to a compact set, since the problem becomes ill-posed and physically unrealistic when perturbations are completely unconstrained.
2.2 Lipschitz Continuity and Lipschitz Regularization
Notwithstanding their impressive performance, modern deep learning systems are sensitive to their inputs: Imperceptibly small but well-chosen perturbations to their inputs can significantly change their outputs (Virmaux & Scaman, 2018; Krishnan et al., 2020). Since modern RL commonly uses deep function approximation for policies as well as value function estimators, this issue is present in the RL setting (Pirotta et al., 2015; Kobayashi, 2022). Lipschitz regularization of deep learning models is one way to improve their robustness against small perturbations (Yang et al., 2020; Jordan & Dimakis, 2020). The Lipschitz constant of a deep model describes its smoothness (Liu et al., 2022; Song et al., 2023). Global Lipschitz constants (and regularization) are concerned with smoothness everywhere in the input space, whereas local Lipschitz properties only require smoothness in certain regions of interest in the input space (e.g. the neighborhoods of training samples) (Jordan & Dimakis, 2020). Therefore, bounding or minimizing the Lipschitz constants, through regularization or other means, will effectively smooth the input-output mapping of the network and increase its robustness against perturbations (Jordan & Dimakis, 2020; Shi et al., 2022). However, if the regularization is too aggressive, it can make the function smoother than necessary at the expense of the primary training objective (Yang et al., 2020). In general, increased robustness against the worst-case typically incurs some reduced performance in the average and best case.
3 Preliminaries
In this section, we introduce the definition of Lipschitz continuity, the robust RL framework, and other terminology and notation used in this paper.
Lipschitz Continuity: Given two metric spaces $(X,d_{X})$ and $(Y,d_{Y})$ with respective distance metrics $d_{X}$ and $d_{Y}$, a function $f:X\to Y$ is $L$-Lipschitz continuous (also called $L$-Lipschitz for short) over $X$ if there exists a constant $L\geq 0$ such that $d_{Y}(f(x_{1}),f(x_{2}))\leq Ld_{X}(x_{1},x_{2})$ for all $x_{1},x_{2}\in X$. The smallest $L=\sup_{x_{1}\neq x_{2}}\frac{d_{Y}(f(x_{1}),f(x_{2}))}{d_{X}(x_{1},x_{2})}$ is the Lipschitz constant of $f$.
Markov Decision Processes (MDPs): An MDP is a tuple $(\mathcal{S},\mathcal{A},T,R,\gamma)$. This paper focuses on MDPs with metric state and action spaces $\mathcal{S}\subseteq\mathbb{R}^{m}$ and $\mathcal{A}\subseteq\mathbb{R}^{n}$ using vector-norm-based distance metrics, such as continuous control tasks. At every time step $t$, the agent takes an action $a_{t}$ to move from the current state $s_{t}$ to a new state $s_{t+1}$. In this paper we assume that the transition is deterministic and specified by the transition function $s_{t+1}=T(s_{t},a_{t})$. The instantaneous reward after each transition is $r_{t}=R(s_{t},a_{t})$ and $\gamma\in(0,1)$ is the discount factor.
Robust MDPs: In robust RL, the MDP formalism is extended to account for perturbations in the transition dynamics. A robust MDP (RMDP) is defined as $(\mathcal{S},\mathcal{A},\mathcal{T},R,\gamma)$, where $\mathcal{S}$, $\mathcal{A}$, $R$, and $\gamma$ are the same as a standard MDP, but $\mathcal{T}$ is an “uncertainty set” of possible transition functions. At every time-step, one particular transition function $T\in\mathcal{T}$ governs the state transition $s_{t+1}=T(s_{t},a_{t})$. The selected transition function $T$ can be random or adversarial. As mentioned earlier, to make the problem well-posed and physically realistic, the perturbations are generally assumed to be bounded. One way to formulate this constraint is to specify a “nominal” transition dynamics $T_{0}$ (e.g., the simulator dynamics) and assume that any perturbed $\hat{T}\in\mathcal{T}$ will always produce a new state within distance $\epsilon$ of what $T_{0}$ would have produced, starting from the same current state and action. Formally, we define the uncertainty set as:
$\mathcal{T}_{\epsilon}=\{\hat{T}\,\big{|}\,\forall(s,a)\in\mathcal{S}\times\mathcal{A}:\,d_{\mathcal{S}}(\hat{T}(s,a),T_{0}(s,a))\leq\epsilon\}$ (1)
The objective is to obtain the optimal policy $\pi^{*}(a|s)$ maximizing the worst-case expected return $\mathcal{J}_{\mathcal{T}_{\epsilon}}(\pi)$ among all possible transition functions: $\mathcal{J}_{\mathcal{T}_{\epsilon}}(\pi)=\inf_{T\in\mathcal{T}_{\epsilon}}\mathbb{E}_{\pi,T}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$. Accordingly, the robust value functions are defined as $V^{\pi}_{\mathcal{T_{\epsilon}}}(s)=\inf_{T\in\mathcal{T}_{\epsilon}}\mathbb{E}_{\pi,T}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|s_{0}=s\right]$ and $Q^{\pi}_{\mathcal{T_{\epsilon}}}(s,a)=\inf_{T\in\mathcal{T}_{\epsilon}}\mathbb{E}_{\pi,T}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|s_{0}=s,a_{0}=a\right]$ (Iyengar, 2005). The RMDP Bellman operator $\mathcal{B}_{\mathcal{T}_{\epsilon}}$ therefore acts by selecting the subsequent state yielding the worst-case state-value:
$\mathcal{B}_{\mathcal{T}_{\epsilon}}Q(s,a)=R(s,a)+\gamma\inf_{T\in\mathcal{T}_{\epsilon}}V(s^{\prime})_{|s^{\prime}=T(s,a)}$ (2)
4 Methodology
4.1 Motivation: State Disturbance in Robust MDP
Based on a strong duality theorem (Blanchet & Murthy, 2019), the RMDP Bellman operator is equivalent to the following optimization problem (Kuang et al., 2022):
$\begin{split}\mathcal{B}_{\mathcal{T}_{\epsilon}}Q(s,a)&=R(s,a)+\gamma\sup_{\lambda\geq 0}\inf_{T\in\mathcal{T}_{\epsilon}}\left(V(s^{\prime})+\lambda(d_{\mathcal{S}}(s^{\prime},s^{\prime}_{0})-\epsilon)\right)_{|s^{\prime}=T(s,a),s^{\prime}_{0}=T_{0}(s,a)}\\ &=R(s,a)+\gamma\inf_{\bar{s}\in B_{\epsilon}(s^{\prime}_{0})}V(\bar{s})\end{split}$ (3)
where $B_{\epsilon}(s^{\prime}_{0})$ is a ball centered at $s^{\prime}_{0}$ with radius $\epsilon$. The advantage of this reformulation, called state disturbance, is that it is more practical to optimize over states than over transition functions.
Eq. 3 still involves an infimum which is difficult to solve exactly. SC-SAC (Kuang et al., 2022) introduces a fast method to approximate the optimum called gradient based regularizer (GBR), using a first-order Taylor expansion and $\infty$-norm for the ball $B_{\epsilon}(s^{\prime}_{0})$:
$\begin{split}\inf_{\bar{s}\in B_{\epsilon}(s^{\prime}_{0})}V(\bar{s})&\approx\inf_{\bar{s}\in B_{\epsilon}(s^{\prime}_{0})}\left(V(s^{\prime}_{0})+\langle\nabla_{s}V(s^{\prime}_{0}),\bar{s}-s^{\prime}_{0}\rangle\right)\\ &=V(s^{\prime}_{0})-\epsilon||\nabla_{s}V(s^{\prime}_{0})||_{1}\end{split}$ (4)
Kuang et al. (2022) observed performance issues with GBR unless $\epsilon$ is rather small. We hypothesize that GBR’s sensitivity to $\epsilon$ is due in part to large and inaccurate estimates of $\nabla_{s}V$ before training is complete. In this case, SC-SAC could be improved by Lipschitz regularization of the critic network, and other modifications that reduce the approximation errors of GBR.
4.2 Lipschitz Continuity of RMDP Value Functions
Lipschitz regularization of the critic network is only justified if the true optimal RMDP value functions are Lipschitz continuous. In this section we show that if $R$ and all $T\in\mathcal{T}_{\epsilon}$ are Lipschitz, so are the optimal value functions. In common continuous control tasks, $R$ and $T$ are Lipschitz almost everywhere except for impulse events. We use a similar proof strategy to Rachelson & Lagoudakis (2010) in their work on Lipschitz MDPs, with some modifications to work with RMDPs.
We first recall a useful lemma for MDPs (Russell & Norvig, 2016):
Lemma 1.
For any two functions $f$ and $g$ on a compact domain, we have
$|\inf_{x}f(x)-\inf_{y}g(y)|$ $\leq\sup_{z}|f(z)-g(z)|$ (5)
$|\sup_{x}f(x)-\sup_{y}g(y)|$ $\leq\sup_{z}|f(z)-g(z)|$ (6)
Proof.
For (5), suppose without loss of generality that $\inf_{x}f(x)\geq\inf_{y}g(y)$ (otherwise, use the following reasoning with $f$ and $g$ swapped). Letting $y^{*}$ denote a minimizer of $g$, we have
$|\inf_{x}f(x)-\inf_{y}g(y)|=\inf_{x}f(x)-g(y^{*})\leq f(y^{*})-g(y^{*})=|f(y^{*})-g(y^{*})|\leq\sup_{z}|f(z)-g(z)|.$ (7)
We know that $y^{*}$ exists since the domain is compact, and we can reintroduce absolute value signs because the the leftmost quantity is non-negative. The reasoning for (6) is similar. ∎
Next we show Lipschitz continuity of the robust optimal value function, as long as $R$ and $T$ are Lipschitz, and $\gamma$ is sufficiently close to 0. Here we assume $T$’s Lipschitz property is expressed using the distance metric $d_{\mathcal{SA}}((s_{1},a_{1}),(s_{2},a_{2}))=d_{\mathcal{S}}(s_{1},s_{2})+d_{\mathcal{A}}(a_{1},a_{2})$.
Proposition 1.
Given any RMDP, suppose that $R$ and all $T\in\mathcal{T}_{\epsilon}$ are Lipschitz with constants $L_{R}$ and $L_{T}$. Let $L_{T^{*}}=\sup_{T\in\mathcal{T}_{\epsilon}}L_{T}$ and suppose $\gamma\in(0,1)$ is selected such that $\gamma L_{T^{*}}<1$. Then the optimal robust value function $V^{*}$ is also Lipschitz, with its Lipschitz constant bounded by
$L_{V^{*}}\leq\frac{L_{R}}{1-\gamma L_{T^{*}}}.$ (8)
Proof.
Following Rachelson & Lagoudakis (2010), we first show Lipschitz continuity for each finite horizon, $n$-step optimal value function (denoted $V_{n}$) by induction on $n$. Then we recover the bound on $L_{V^{*}}$ in the limit $n\rightarrow\infty$. In the base case $n=0$, we have $V_{0}(s)=\max_{a}R(s,a)$, and
$|V_{0}(s_{1})-V_{0}(s_{2})|=|\max_{a_{1}}R(s_{1},a_{1})-\max_{a_{2}}R(s_{2},a_{2})|\leq\max_{a}|R(s_{1},a)-R(s_{2},a)|\leq L_{R}d_{\mathcal{S}}(s_{1},s_{2}),$
where the first inequality uses Lemma 1 and the second uses $d_{\mathcal{SA}}((s_{1},a),(s_{2},a))=d_{\mathcal{S}}(s_{1},s_{2})$.
For the inductive case, we note that the $(n+1)$-step value function satisfies
$V_{n+1}(s)=\max_{a}\left(R(s,a)+\gamma\inf_{T\in\mathcal{T}_{\epsilon}}V_{n}(T(s,a))\right).$ (9)
Assuming the inductive hypothesis that $V_{n}$ is $L_{n}$-Lipschitz continuous, we derive $L_{n+1}$ as follows:
$|V_{n+1}(s_{1})-V_{n+1}(s_{2})|$ (10)
$=$ $\left|\max_{a_{1}}\left(R(s_{1},a_{1})+\gamma\inf_{T_{1}}V_{n}(T_{1}(s_{1},a_{1}))\right)-\max_{a_{2}}\left(R(s_{2},a_{2})+\gamma\inf_{T_{2}}V_{n}(T_{2}(s_{2},a_{2}))\right)\right|$ (11)
$\leq$ $\max_{a}\left|\left(R(s_{1},a)+\gamma\inf_{T_{1}}V_{n}(T_{1}(s_{1},a))\right)-\left(R(s_{2},a)+\gamma\inf_{T_{2}}V_{n}(T_{2}(s_{2},a))\right)\right|$ (12)
$\leq$ $\left|R(s_{1},a^{*})-R(s_{2},a^{*})\right|+\gamma\left|\inf_{T_{1}}V_{n}(T_{1}(s_{1},a^{*}))-\inf_{T_{2}}V_{n}(T_{2}(s_{2},a^{*}))\right|$ (13)
$\leq$ $L_{R}d_{\mathcal{S}}(s_{1},s_{2})+\gamma\sup_{T}\left|V_{n}(T(s_{1},a^{*}))-V_{n}(T(s_{2},a^{*}))\right|$ (14)
$=$ $L_{R}d_{\mathcal{S}}(s_{1},s_{2})+\gamma\left|V_{n}(T^{*}(s_{1},a^{*}))-V_{n}(T^{*}(s_{2},a^{*}))\right|\leq(L_{R}+\gamma L_{n}L_{T^{*}})d_{\mathcal{S}}(s_{1},s_{2}),$ (15)
where $a^{*}$ and $T^{*}$ are the respective maximizers. We used Lemma 1 in (12) and (14) and simplified $d_{\mathcal{SA}}$ similarly to the base case. Expanding the recurrence $L_{n+1}\leq L_{R}+\gamma L_{n}L_{T^{*}}$, taking the limit $n\rightarrow\infty$ and using properties of geometric series, the claim follows. ∎
Proposition 1 justifies Lipschitz regularization of the critic network as a means of conditioning $\nabla_{s}V$ in GBR. However, $\nabla_{s}V$ is evaluated based on its relation to $Q$, which is the function actually approximated by the critic network. So we proceed to show first that $Q^{*}$ is also Lipschitz, and second that its constant $L_{Q^{*}}$ upper bounds $L_{V^{*}}$.
Proposition 2.
Given a Lipschitz RMDP as in Proposition 1, $Q^{*}$ is also Lipschitz continuous with the same bound on $L_{Q^{*}}$.
Proof.
$Q^{*}$ and $V^{*}$ are related by
$Q^{*}(s,a)=R(s,a)+\gamma\inf_{T\in\mathcal{T}_{\epsilon}}V^{*}(T(s,a)).$ (16)
Therefore we have
$|Q^{*}(s_{1},a_{1})-Q^{*}(s_{2},a_{2})|$ $\leq|R(s_{1},a_{1})-R(s_{2},a_{2})|+\gamma\sup_{T}|V^{*}(T(s_{1},a_{1}))-V^{*}(T(s_{2},a_{2}))|$ (17)
$\leq(L_{R}+\gamma L_{V^{*}}L_{T^{*}})\cdot d_{\mathcal{SA}}((s_{1},a_{1}),(s_{2},a_{2}))$ (18)
$\leq(L_{R}+\gamma\frac{L_{R}}{1-\gamma L_{T^{*}}}L_{T^{*}})\cdot d_{\mathcal{SA}}((s_{1},a_{1}),(s_{2},a_{2}))$ (19)
$=\frac{L_{R}}{1-\gamma L_{T^{*}}}\cdot d_{\mathcal{SA}}((s_{1},a_{1}),(s_{2},a_{2})),$ (20)
where (17) uses Lemma 1 yet again and (19) uses the bound on $L_{V^{*}}$ from Proposition 1. Therefore $L_{Q^{*}}\leq L_{R}/(1-\gamma L_{T^{*}})$. ∎
Finally, we justify regularization of the critic approximation of $Q^{*}$ as a proxy for $V^{*}$ by proving that $L_{Q^{*}}$ is an upper bound on $L_{V^{*}}$:
Proposition 3.
Given a Lipschitz RMDP as in Propositions 1 and 2, $L_{V^{*}}\leq L_{Q^{*}}$.
Proof.
$V^{*}$ and $Q^{*}$ are also related by $V^{*}(s)=\max_{a}Q^{*}(s,a)$. Therefore we have
$|V^{*}(s_{1})-V^{*}(s_{2})|$ $=|\max_{a_{1}}Q^{*}(s_{1},a_{1})-\max_{a_{2}}Q^{*}(s_{2},a_{2})|$ (21)
$\leq\max_{a}|Q^{*}(s_{1},a)-Q^{*}(s_{2},a)|$ (22)
$=|Q^{*}(s_{1},a^{*})-Q^{*}(s_{2},a^{*})|$ (23)
$\leq L_{Q^{*}}d_{\mathcal{S}}(s_{1},s_{2}),$ (24)
where $a^{*}$ is again the relevant maximizer. Once more we have used Lemma 1 in (22), and simplified $d_{\mathcal{SA}}$ in (24) similarly to the base case of Proposition 1. Since $s_{1}$ and $s_{2}$ are arbitrary, $|V^{*}(s_{1})-V^{*}(s_{2})|/d_{\mathcal{S}}(s_{1},s_{2})\leq L_{Q^{*}}$ for all $s_{1}\neq s_{2}$. Therefore the Lipschitz constant $L_{V^{*}}\leq L_{Q^{*}}$ also. ∎
4.3 Explicit Lipschitz Value Estimation (ELVEn)
In this section, we introduce Explicit Lipschitz Value Estimation (ELVEn), which combines Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2014) and Lipschitz regularization to reduce the sensitivity of GBR to $\epsilon$. GBR is closely related to FGSM since
$V(s^{\prime}_{0})-\epsilon||\nabla_{s}V(s^{\prime}_{0})||_{1}=V(s^{\prime}_{0})+\langle\nabla_{s}V(s^{\prime}_{0}),\hat{s}-s^{\prime}_{0}\rangle\quad\text{when}\quad\hat{s}=s^{\prime}_{0}-\epsilon\cdot\text{sign}(\nabla_{s}V(s^{\prime}_{0})).$ (25)
This adversarial computation of $\hat{s}$ is precisely FGSM, which we denote by $\hat{s}=Adv(s,\epsilon,V)$. This means that GBR is equivalent to evaluating the first-order Taylor expansion of $V$ at $Adv(s,\epsilon,V)$. However, GBR does not explicitly compute $Adv(s,\epsilon,V)$ or reevaluate $V$; it only explicitly calculates $\epsilon\cdot\nabla_{s}V(s^{\prime}_{0})$. We hypothesize that explicitly evaluating $V$ at $Adv(s,\epsilon,V)$, rather than implicitly evaluating its first-order Taylor expansion there, will provide a more accurate estimate of $\inf_{\bar{s}\in B_{\epsilon}(s^{\prime}_{0})}V(\bar{s})$. This does incur the cost of one additional forward-pass evaluation of $V$ per update, but has the potential benefit of reducing sensitivity to $\epsilon$ and improving robustness.
The foregoing modification still does not mitigate the issue of a large and ill-conditioned $\nabla_{s}V$. To also mitigate this issue, based on Propositions 1-3, we perform Lipschitz regularization on the critic network. However, optimizing the training loss subject to a hard Lipschitz constraint is an infeasible problem. It is not only intractable to determine the precise Lipschitz constants in advance, but also computing the Lipschitz constant of neural networks is considered NP-hard (Virmaux & Scaman, 2018; Kobayashi, 2022). Instead of solving an intractable constrained optimization, we add a penalty term to the objective that is proportional to the square of
$\frac{|Q(s,a)-Q(\hat{s},a)|}{d_{\mathcal{SA}}((s,a),(\hat{s},a))}=\frac{|Q(s,a)-Q(\hat{s},a)|}{\epsilon}.$ (26)
4.4 ELVEn-SAC
We apply ELVEn as proposed above to Soft Actor-Critic (SAC) (Haarnoja et al., 2018), a popular off-policy actor-critic algorithm, and call the resulting implementation “ELVEn-SAC.” The implementation is described below and with pseudocode in Algorithm 1. We apply the reparameterization trick to the policy network $\pi_{\psi}(a|s)$ when predicting the action distribution. We use two separate critic networks $Q_{\theta_{i}}(s,a),i=\{1,2\}$ to perform clipped double Q-learning (Fujimoto et al., 2018) and use two target networks $Q_{\hat{\theta}_{i}}(s,a)$ to soft-update the corresponding critic. The SAC temperature $\alpha$ in ELVEn-SAC is a trainable parameter responsible for adjusting the weight of the entropy term.
The critic loss of ELVEn-SAC consists of two terms: a prediction loss and a Lipschitz regularization loss. Given a transition data-point $(s,a,s^{\prime},r)$ and perturbation radius $\epsilon$, the learning target $\hat{Q}$ is
$\hat{Q}(s,a)=r+\gamma\mathbb{E}_{a^{\prime}\sim\pi_{\psi}(\cdot|s^{\prime})}\left[\min_{i=1,2}Q_{\hat{\theta}_{i}}(Adv(s^{\prime},\epsilon,Q_{\hat{\theta}_{i}}),a^{\prime})-\alpha\log\pi_{\psi}(a^{\prime}|s^{\prime})\right].$ (27)
The prediction loss minimizes the residual in the robust Bellman operator by learning $\hat{Q}(s,a)$:
$\mathcal{L}^{\text{Pred}}_{Q}(\theta)=\left(Q_{\theta}(s,a)-\hat{Q}(s,a)\right)^{2}$ (28)
The Lipschitz regularization loss minimizes the critic network change at adversarial states:
$\mathcal{L}^{\text{Lips}}_{Q}(\theta)=\left(Q_{\theta}(s,a)-Q_{\theta}(Adv(s,\epsilon,Q_{\theta}),a)\right)^{2}$ (29)
The final critic loss is then defined as following, where $\lambda$ is a hyperparameter controlling the importance of regularization versus temporal difference learning:
$\mathcal{L}_{Q}(\theta)=\mathcal{L}^{\text{Pred}}_{Q}(\theta)+\lambda\mathcal{L}^{\text{Lips}}_{Q}(\theta)$ (30)
The actor loss maximizes the entropy of the policy, relative to the current action value estimates:
$\mathcal{L}_{\pi}(\psi)=\mathbb{E}_{\bar{a}\sim\pi_{\psi}(\cdot|s)}\left[\alpha\log\pi_{\psi}(\bar{a}|s)-\min_{i=1,2}Q_{\theta_{i}}(Adv(s,\epsilon,Q_{\theta_{i}}),\bar{a})\right]$ (31)
We tune the temperature parameter $\alpha$ with the following loss, where $\mathcal{H}$ is a constant target entropy:
$\mathcal{L}_{\alpha}(\alpha)=-\alpha\cdot\mathbb{E}_{\bar{a}\sim\pi_{\psi}(\cdot|s)}\left[\log\pi_{\psi}(\bar{a}|s)+\mathcal{H}\right]$ (32)
Algorithm 1 Explicit Lipschitz Value Estimation SAC (ELVEn-SAC)
Critics $Q_{\theta_{1}}(s,a)$, $Q_{\theta_{2}}(s,a)$ and targets $Q_{\hat{\theta}_{1}}(s,a)$, $Q_{\hat{\theta}_{2}}(s,a)$. Policy $\pi_{\psi}(a|s)$. Temperature parameter $\alpha$. Learning rates $\eta_{\theta}$, $\eta_{\psi}$, $\eta_{\alpha}$. Soft-update coefficient $\tau$. Replay buffer $\mathcal{D}$.
Update $Q_{\theta_{1}}(s,a)$, $Q_{\theta_{2}}(s,a)$ and $\pi_{\psi}(a|s)$.
for each time step $t$ do
     $a_{t}\sim\pi_{\psi}(\cdot|s_{t})$, $s_{t+1}=T(s_{t},a_{t})$, $r_{t}=R(s_{t},a_{t})$
     $\mathcal{D}\leftarrow\mathcal{D}\cup(s_{t},a_{t},s_{t+1},r_{t})$
     for each training step do
         $(s,a,s^{\prime},r)\sim\mathcal{D}$
         $\theta_{i}\leftarrow\theta_{i}-\eta_{\theta}\nabla_{\theta_{i}}\mathcal{L}_{Q}(\theta_{i})$, $i\in\{1,2\}$
         $\psi\leftarrow\psi-\eta_{\psi}\nabla_{\psi}\mathcal{L}_{\pi}(\psi)$
         $\alpha\leftarrow\alpha-\eta_{\alpha}\nabla_{\alpha}\mathcal{L}_{\alpha}(\alpha)$
         $\hat{\theta}_{i}\leftarrow(1-\tau)\hat{\theta}_{i}+\tau\theta_{i}$, $i\in\{1,2\}$
     end for
end for
5 Experiments
5.1 Experimental Setup
We conduct experiments on four MuJoCo (Todorov et al., 2012) environments: Ant, HalfCheetah, Hopper and Walker2d. We firstly discuss our Lipschitz assumption on these environments. The evolution of a Lipschitz continuous dynamical system is not subject to discontinuous changes. In test environments, discontinuity only comes from the impulses that the agent hits the floor. The Lipschitz condition of transition dynamics is locally satisfied in most of the state-action space, except in the regions where the impulses happen. For the reward formulation, all four environments have a consistent form $R(s,a)=\alpha_{1}v+\alpha_{2}-\alpha_{3}\sum_{i}a^{2}_{i}$, where $v$ is the forward velocity, $a_{i}$ is the $i$-th element of the action vector $a$, and $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{3}$ are task-specific non-negative constants. This reward function is Lipschitz thanks to two facts: (1) $v$ is included in the state vector $s$, and (2) the range of actions are $[-1,1]$, meaning that the gradient norm of the smooth function $\sum_{i}a^{2}_{i}$ is upper bounded.
To compare with ELVEn-SAC, we implement two baseline algorithms: SAC (Haarnoja et al., 2018) and SC-SAC (Kuang et al., 2022). For each task, we train the policy on the nominal environment for 1M steps, run 4 repetitions with different random seeds, and evaluate the average performance across policies at multiple time-points near the end of training: 970K, 980K, 990K and 1M steps. This is a total of $4\times 4=16$ evaluations per task. Each evaluation uses 10 independent episodes with the current policy. Table 1 lists the hyperparameter values held fixed in our experiments.
Parameter Value
Optimizer Adam
Hidden layer size [256, 256]
Learning rate of network $3\times 10^{-4}$
Learning rate of $\alpha$ $3\times 10^{-4}$
Discount factor $\gamma$ 0.99
Soft-update coefficient $\tau$ 0.005
Parameter Value
Reward scale 1.0
Target entropy $-\dim(\mathcal{A})$
Exploration steps $1\times 10^{4}$
Mini-batch size 256
Replay buffer capacity $1\times 10^{6}$
Update frequency 1 update per step
Table 1: Shared hyperparameters for SAC, SC-SAC and ELVEn-SAC. Left: Hyperparameters for MLP actor and critic network. Right: Hyperparameters for policy training.
5.2 Empirical Performance and Robustness
Figure 1: Relative performance and robustness. We set $\epsilon=0.005$ for SC-SAC and ELVEn-SAC, and $\lambda=0.1$ for ELVEn-SAC. Red columns show average reward across 160 samples for each algorithm, task, and environment, and darker red represents higher reward. Blue-yellow columns with title format “$A_{1}<A_{2}$” show the robustness indicator $\mathbf{1}_{T,A_{1},A_{2}}$ (see text) at the “pixel” corresponding to environment $T$, with dark blue = $-1$, yellow = $+1$, and teal = 00. That is, the inequality is satisfied at yellow pixels, and its reverse is satisfied at dark-blue pixels. There is no statistically significant difference at teal pixels.
[…]
$t_{T,A_{1},A_{2}}=\frac{m_{1}-m_{2}}{\sqrt{z^{2}_{1}+z^{2}_{2}}},$ (33)
where $m_{i}$ and $z_{i}$ are the sample mean and standard error in net reward for algorithm $A_{i}$, over its size-160 sample. The alternative hypothesis was that $A_{1}$ has lower mean reward than $A_{2}$, which we denote $A_{1}<A_{2}$. This hypothesis test was conducted for each direction $A_{1}<A_{2}$ and $A_{2}<A_{1}$, on each pairing of two of the three algorithms, and each of the $11\times 1$ 11 perturbed environments across the $4$ tasks, for a total of $11\times 11\times 2\times 3\times 4=2904$ hypothesis tests. Since a large number of hypothesis tests runs the risk of many false positives, we used Bonferroni correction on a p-value of 0.05, meaning we only accepted the alternative hypothesis for p-values less than $0.05/2904$.
For each environment $T$ and algorithm pair $\{A_{1},A_{2}\}$, we defined an indicator variable $\mathbf{1}_{T,A_{1},A_{2}}$ to indicate whether we accept alternative hypothesis $A_{1}<A_{2}$, $A_{2}<A_{1}$, or neither:
$\mathbf{1}_{T,A_{1},A_{2}}=\begin{cases}+1&\text{ if }A_{1}<A_{2}\\ -1&\text{ if }A_{1}>A_{2}\\ 0&\text{ otherwise }\\ \end{cases}$ (34)
This performance indicator is visualized in Figure 1 for each task and pair of algorithms.
We then define the relative “robustness” of one algorithm vs. another by averaging this variable across all environments in the task:
$\text{robustness}(\mathcal{T}_{\epsilon},A_{1},A_{2})=\frac{1}{11\times 11}\sum_{T\in\mathcal{T}_{\epsilon}}\mathbf{1}_{T,A_{1},A_{2}}$ (35)
Values near 1 signify that $A_{2}$ is consistently, statistically significantly more performant than $A_{1}$ in perturbed environments.
In robust RL, improved robustness often comes at the cost of reduced performance on the default environment $T_{0}$. We also quantify this performance change using the t-statistic $t_{T_{0},A_{1},A_{2}}$ on the default environment $T_{0}$. This difference of mean reward is already normalized by the standard error, so its value relative to ±$\pm 1$ indicates the statistical significance of the change in performance.
Both “robustness” and this t-statistic are shown in Table 2, for each of SC-SAC and ELVEn-SAC relative to SAC, on all four tasks. Although the results vary by task, we can observe that by both metrics, ELVEn-SAC outperforms SC-SAC on a majority of tasks. Statistically speaking, ELVEn-SAC’s performance on $T_{0}$ is never much worse, and often better, than SAC.
t-statistic robustness
Task SC-SAC ELVEn-SAC SC-SAC ELVEn-SAC
Ant -1.410 -0.039 -0.215 0.017
HalfCheetah 2.462 0.320 0.107 -0.033
Walker2d 6.203 10.321 0.331 0.818
Hopper -1.073 6.240 0.140 0.628
[…]
Figure 2: The sensitivity of SC-SAC and ELVEn-SAC to $\epsilon$ on Hopper. We compare the performance when $\epsilon\in\{0.001,0.003,0.005,0.007,0.01\}$. In ELVEn-SAC, $\lambda$ is fixed to 0.1.
Figure 3: The sensitivity of ELVEn-SAC to $\lambda$ on Hopper when $\epsilon\in\{0.007,0.01\}$. We attach the performance of SC-SAC with same $\epsilon$ for comparison.
5.3 Hyper-parameter Analyses
We empirically analyze the sensitivity of SC-SAC and ELVEn-SAC to $\epsilon$. Figure 2 shows the evaluation results of policies trained with different $\epsilon$. When $\epsilon$ is within a specific range, both SC-SAC and ELVEn-SAC achieve good generalization in perturbed environments. For larger $\epsilon$, SC-SAC’s performance drops while ELVEn-SAC’s performance is more stable. Therefore, ELVEn-SAC is less sensitive to the choice of $\epsilon$ than SC-SAC, showing robustness against larger perturbations.
Figure 3 shows how robustness of ELVEn-SAC can improve and eventually exceed that of SC-SAC as $\lambda$ increases. We also inspect the effect of $\lambda$ by visualizing the smoothness of the regularized and unregularized $Q$ functions in Figure 4. The figure confirms that $Q$ becomes smoother with more aggressive regularization (higher $\lambda$), and is comparable for SC-SAC and ELVEn-SAC without regularization ($\lambda=0$).
Performance plots for the full range of $\epsilon$ and $\lambda$ tested are included in the Appendix.
(a) SC-SAC
(b) $\lambda=0$
(c) $\lambda=0.1$
(d) $\lambda=0.5$
Figure 4: The 3D landscape of predicted $Q(s,a)$ on Hopper for SC-SAC (red) and ELVEn-SAC (green). The variable $s[0]$ and $s[1]$ denote the first two dimensions of state space. We vary $s[0]$ and $s[1]$, and fix the other states and actions. We keep $\epsilon=0.005$ for both SC-SAC and ELVEn-SAC. Note the different y-axis scales.
6 Conclusion
In this paper, we tested our hypotheses that Lipschitz regularization of the value function can enhance robustness to environmental perturbations. We introduced Explicit Lipschitz Value Estimation (ELVEn), a combination of Fast Gradient Sign Method and Lipschitz regularization, and integrated it into an off-policy algorithm ELVEn-SAC. Our experimental results show that ELVEn-SAC is less sensitive to perturbations than prior work. In future work, we aim to further explore ELVEn in conjunction with local Lipschitz regularization, evaluate ELVEn on a wider range of neural architectures and tasks, and test its utility for sim-to-real transfer on real physical robotic hardware.
Appendix
A. Details of Implementation and Experiments
We follow Kuang et al. (2022) to implement observation normalization in SAC, SC-SAC and ELVEn-SAC. For each task, we firstly run SAC without observation normalization for 1M steps, computing the mean and standard deviation with the observations in the replay buffer. Then we use these fixed values to normalize the observations.
We run all experiments on a shared GPU cluster where the primary GPU model is Quadro RTX 6000. We estimate the computation cost of three algorithms in Table 3. However, due to the resource allocation on the shared cluster, making an objective comparison can be challenging.
Task (Unit: s) Ant HalfCheetah Hopper Walker2d
SAC 99.95 100.31 98.32 107.05
SC-SAC 177.37 212.15 198.18 194.33
ELVEn-SAC 160.33 158.75 155.40 156.39
Table 3: The total time cost of updating 9900 batches in SAC, SC-SAC and ELVEn-SAC.
The complete performance plots of SAC, SC-SAC and ELVEn-SAC are following, where $\epsilon\in\{0.001,0.003,0.005,0.007,0.01\}$ and $\lambda\in\{0,0.1,0.2,0.5\}$.