### v2

# Explicit Lipschitz Value Estimation Enhances Policy Robustness Against Perturbation

Abstract
In robotic control tasks, policies trained by reinforcement learning (RL) in simulation often experience a performance drop when deployed on physical hardware, due to modeling error, measurement error, and unpredictable perturbations in the real world. Robust RL methods account for this issue by approximating a worst-case value function during training, but they can be sensitive to approximation errors in the value function and its gradient before training is complete. In this paper, we hypothesize that Lipschitz regularization can help condition the approximated value function gradients, leading to improved robustness after training. We test this hypothesis by combining Lipschitz regularization with the Fast Gradient Sign Method to reduce approximation errors when evaluating the value function under adversarial perturbations. Our empirical results demonstrate the benefits of this approach over prior work on several continuous control benchmarks.
1 Introduction
Recently, reinforcement learning (RL) has demonstrated impressive capabilities in robotics control tasks. In practice, RL policies are often trained in physics-based simulators before being transferred to the real world. However, the discrepancy between the simulation and reality, known as the reality gap, can hamper performance when simulated policies are deployed in the real world. Common discrepancies include system identification errors (incorrect mass, friction coefficients, etc. in simulation), idealized physical modeling (e.g., perfectly rigid bodies), and state estimation errors (e.g. sensor noise) in the real world. These discrepancies can be viewed as small differences in the transition dynamics of the simulator vs. those of the real world [1].
Two approaches to bridge the reality gap are to improve the fidelity of the simulators, and to improve the robustness of RL-trained policies against perturbations. While some works focus on the first approach, building simulators that more faithfully replicate real-world physics [2], a widely used technique in the second approach is domain randomization (DR) [3, 4, 5]. DR works with existing simulators by randomizing the physical parameters (e.g., mass and friction), which effectively exposes the policies to diverse environments during training.
Robust RL uses adversarial perturbations rather than randomly sampled perturbations during training [6]. This promotes acceptable performance even in a worst case scenario where reality differs from the simulated environment in a pathological way. The trade-off is that best- or average-case performance may be slightly worse compared to policies trained with DR, but this is a small price to pay in high-risk deployment scenarios where dangerous operating modes must be avoided at all costs. Robust RL requires a sub-routine that can approximate the worst-case perturbation at a given state. These approximations are often made using the current value function estimate [7], in which case the training process is sensitive to errors in the value function estimates before learning has converged.
Lipschitz regularization is an effective strategy for reducing sensitivity of neural networks to adversarial examples. It has been used in both computer vision [8] and in RL, to smooth the policy network actions [9] as well as critic network outputs [10]. However, to our knowledge, existing work has not yet studied the efficacy of Lipschitz regularization for reducing sensitivity to value estimates during robust RL training. In this paper, we show that Lipschitz regularization can improve the robustness of the trained policies in a number of continuous control benchmarks. When generating adversarial perturbations during training, we incorporate a novel use of the Fast Gradient Sign Method (FGSM) [11] to further reduce errors in value function estimates. Unlike past work that uses FGSM after training [12] or implicitly as part of a Taylor approximation [7], we re-evaluate the value function on the explicitly computed adversarial perturbation to remove another source of approximation error.
2 Related Work
Robust RL
With the advent of adversarial examples and emerging AI safety concerns, robust RL [6] has drawn increasing attention recently. Adversarial examples can be incorporated after [12] or during [13, 1, 14, 15] RL training to improve robustness against worst-case transition dynamics in some compact set of possible environments. Robust RL is relevant to the reality gap as well as other unforeseen circumstances that may arise after deployment. These include unintentional accidents in safety-critical application domains such as power grids and autonomous vehicles [16], and intentional deceptions launched by malicious actors [17, 18, 19, 20]. In the latter case, the robust RL optimization is treated as a zero-sum game between the agent and an adversarial opponent [18, 19]. Usually, the size of adversarial perturbations is constrained to a compact set, since the problem becomes ill-posed and physically unrealistic when perturbations are completely unconstrained.
Lipschitz Continuity and Regularization
Notwithstanding their impressive performance, modern deep learning systems are sensitive to their inputs: Imperceptibly small but well-chosen input perturbations can significantly change their outputs [21, 22]. Since modern RL commonly uses deep function approximation for policies and value functions, this issue is present in the RL setting [23, 10]. Lipschitz regularization of deep learning models is one way to improve their robustness against small perturbations [24, 25]. The Lipschitz constant of a deep model describes its smoothness [26, 9]. Global Lipschitz constants (and regularization) are concerned with smoothness everywhere in the input space, whereas local Lipschitz properties only require smoothness in certain regions of interest in the input space (e.g. the neighborhoods of training samples) [25]. Therefore, bounding or minimizing the Lipschitz constants, through regularization or other means, will effectively smooth the input-output mapping of the network and increase its robustness against perturbations [25, 27]. However, if the regularization is too aggressive, it can make the function smoother than necessary at the expense of the primary training objective [24]. In general, increased robustness against the worst-case typically incurs some reduced performance in the average and best case.
3 Preliminaries
In this section, we introduce the definition of Lipschitz continuity, the robust RL framework, and other terminology and notation used in this paper.
Lipschitz Continuity
Given two metric spaces $(X,d_{X})$ and $(Y,d_{Y})$ with respective distance metrics $d_{X}$ and $d_{Y}$, a function $f:X\to Y$ is $L$-Lipschitz continuous (also called $L$-Lipschitz for short) over $X$ if there exists a constant $L\geq 0$ such that $d_{Y}(f(x_{1}),f(x_{2}))\leq Ld_{X}(x_{1},x_{2})$ for all $x_{1},x_{2}\in X$. The smallest $L=\sup_{x_{1}\neq x_{2}}\frac{d_{Y}(f(x_{1}),f(x_{2}))}{d_{X}(x_{1},x_{2})}$ is the Lipschitz constant of $f$.
Markov Decision Processes (MDPs)
An MDP is a tuple $(\mathcal{S},\mathcal{A},T,R,\gamma)$. This paper focuses on MDPs with metric state and action spaces $\mathcal{S}\subseteq\mathbb{R}^{m}$ and $\mathcal{A}\subseteq\mathbb{R}^{n}$ using vector-norm-based distance metrics $d_{\mathcal{S}}(\cdot,\cdot)$ and $d_{\mathcal{A}}(\cdot,\cdot)$ respectively, such as continuous control tasks. At every time step $t$, the agent takes an action $a_{t}$ to move from the current state $s_{t}$ to a new state $s_{t+1}$. In this paper we assume that the transition is deterministic and specified by the transition function $s_{t+1}=T(s_{t},a_{t})$. The instantaneous reward after each transition is $r_{t}=R(s_{t},a_{t})$ and $\gamma\in(0,1)$ is the discount factor.
Robust MDPs:
In robust RL, the MDP formalism is extended to account for perturbations in the transition dynamics. A robust MDP (RMDP) is defined as $(\mathcal{S},\mathcal{A},\mathcal{T},R,\gamma)$, where $\mathcal{S}$, $\mathcal{A}$, $R$, and $\gamma$ are the same as a standard MDP, but $\mathcal{T}$ is an “uncertainty set” of possible transition functions. At every time-step, one particular transition function $T\in\mathcal{T}$ governs the state transition $s_{t+1}=T(s_{t},a_{t})$ and $T$ can be random or adversarial. The objective of RMDP is to obtain the optimal policy $\pi^{*}(a|s)$ maximizing the worst-case expected return $\mathcal{J}_{\mathcal{T}}(\pi)$ across all possible transition functions: $\mathcal{J}_{\mathcal{T}}(\pi)=\inf_{T\in\mathcal{T}}\mathbb{E}_{\pi,T}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$. Accordingly, the value functions of RMDP are defined as $V^{\pi}_{\mathcal{T}}(s)=\inf_{T\in\mathcal{T}}\mathbb{E}_{\pi,T}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|s_{0}=s\right]$ and $Q^{\pi}_{\mathcal{T}}(s,a)=\inf_{T\in\mathcal{T}}\mathbb{E}_{\pi,T}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|s_{0}=s,a_{0}=a\right]$ [6].
4 Methodology
4.1 Motivation: State Disturbance in Robust MDP
To make the problem well-posed and physically realistic, the uncertainty set $\mathcal{T}$ is generally assumed to be bounded. One way to formulate this constraint is called state disturbance [7], which specifies a “nominal” transition function $T_{0}$ (e.g., the simulator dynamics) and assumes that any perturbed $\hat{T}\in\mathcal{T}$ will always produce a new state within distance $\epsilon$ of what $T_{0}$ would have produced, starting from the same current state and action. Formally, we define the uncertainty set as
$\mathcal{T}_{\epsilon}=\{\hat{T}\,\big{|}\,\forall(s,a)\in\mathcal{S}\times\mathcal{A}:\,d_{\mathcal{S}}(\hat{T}(s,a),T_{0}(s,a))\leq\epsilon\}$ (1)
The robust Bellman operator $\mathcal{B}_{\mathcal{T}_{\epsilon}}Q$ therefore acts by selecting the subsequent state yielding the worst-case state-value as Eq. 2, and is equivalent to the following optimization problem [7] according to a strong duality theorem [28]:
$\mathcal{B}_{\mathcal{T}_{\epsilon}}Q(s,a)$ $=R(s,a)+\gamma\inf_{T\in\mathcal{T}_{\epsilon}}V(s^{\prime})_{|s^{\prime}=T(s,a)}$ (2)
$=R(s,a)+\gamma\sup_{\lambda\geq 0}\inf_{T\in\mathcal{T}_{\epsilon}}\left(V(s^{\prime})+\xi(d_{\mathcal{S}}(s^{\prime},s^{\prime}_{0})-\epsilon)\right)_{|s^{\prime}=T(s,a),s^{\prime}_{0}=T_{0}(s,a)}$ (3)
$=R(s,a)+\gamma\inf_{\bar{s}\in B_{\epsilon}(s^{\prime}_{0})}V(\bar{s})$ (4)
where $B_{\epsilon}(s^{\prime}_{0})$ is a ball centered at $s^{\prime}_{0}$ with radius $\epsilon$. The advantage of this reformulation is that it is more practical to optimize over states than over transition functions. However, Eq. 4 still involves an infimum which is difficult to solve exactly. To address this issue, [7] proposed State-Conservative SAC (SC-SAC) by introducing a fast method to approximate the optimum, called gradient based regularizer (GBR), with a first-order Taylor expansion and $\infty$-norm for the ball $B_{\epsilon}(s^{\prime}_{0})$:
$\begin{split}\inf_{\bar{s}\in B_{\epsilon}(s^{\prime}_{0})}V(\bar{s})&\approx\inf_{\bar{s}\in B_{\epsilon}(s^{\prime}_{0})}\left(V(s^{\prime}_{0})+\langle\nabla_{s}V(s^{\prime}_{0}),\bar{s}-s^{\prime}_{0}\rangle\right)\\ &=V(s^{\prime}_{0})-\epsilon||\nabla_{s}V(s^{\prime}_{0})||_{1}\end{split}$ (5)
[7] observed performance issues with GBR unless $\epsilon$ is rather small. We hypothesize that GBR’s sensitivity to $\epsilon$ is due in part to large and inaccurate estimates of $\nabla_{s}V$ before training is complete. In this case, SC-SAC could be improved by Lipschitz regularization of the critic network, and other modifications that reduce the approximation errors of GBR.
4.2 Lipschitz Continuity of RMDP Value Functions
Lipschitz regularization of the critic network is only justified if the true optimal RMDP value functions are Lipschitz continuous. In this section we present theoretical results showing that if $R$ and all $T\in\mathcal{T}_{\epsilon}$ are Lipschitz, so are the optimal value functions. In common continuous control tasks, $R$ can be designed to be Lipschitz on the state-action space (e.g. quadratic functions on compact domains), and Lipschitz continuity is satisfied in some nonlinear system transition dynamics [29]. We use a similar proof strategy to [30] in their work on Lipschitz MDPs, with some modifications to work with RMDPs.
We first recall a useful lemma for MDPs [31]:
Lemma 1.
For any two continuous functions $f$ and $g$ on a compact domain, we have
$|\inf_{x}f(x)-\inf_{y}g(y)|$ $\leq\sup_{z}|f(z)-g(z)|$ (6)
$|\sup_{x}f(x)-\sup_{y}g(y)|$ $\leq\sup_{z}|f(z)-g(z)|$ (7)
Proof.
See Appendix A.1.
Next we confirm Lipschitz continuity of the robust optimal value function, as long as $R$ and $T$ are Lipschitz, and $\gamma$ is sufficiently close to 0. Here we assume $T$’s Lipschitz property is expressed using the distance metric $d_{\mathcal{SA}}((s_{1},a_{1}),(s_{2},a_{2}))=d_{\mathcal{S}}(s_{1},s_{2})+d_{\mathcal{A}}(a_{1},a_{2})$.
Proposition 1.
Given any RMDP, suppose that $R$ and all $T\in\mathcal{T}_{\epsilon}$ are Lipschitz with constants $L_{R}$ and $L_{T}$. Let $L_{T^{*}}=\sup_{T\in\mathcal{T}_{\epsilon}}L_{T}$ and suppose $\gamma\in(0,1)$ is selected such that $\gamma L_{T^{*}}<1$. Then the optimal robust value function $V^{*}$ is also Lipschitz, with its Lipschitz constant bounded by
$L_{V^{*}}\leq\frac{L_{R}}{1-\gamma L_{T^{*}}}.$ (8)
Proof.
See Appendix A.2.
Proposition 1 justifies Lipschitz regularization of the critic network as a means of conditioning $\nabla_{s}V$ in GBR. However, $\nabla_{s}V$ is evaluated based on its relation to $Q$, which is the function actually approximated by the critic network. So we proceed to show first that $Q^{*}$ is also Lipschitz, and second that its constant $L_{Q^{*}}$ upper bounds $L_{V^{*}}$.
Proposition 2.
Given a Lipschitz RMDP as in Proposition 1, $Q^{*}$ is also Lipschitz continuous with the same bound on $L_{Q^{*}}$, meaning
$L_{Q^{*}}\leq\frac{L_{R}}{1-\gamma L_{T^{*}}}.$ (9)
Proof.
See Appendix A.3.
Finally, we justify regularization of the critic approximation of $Q^{*}$ as a proxy for $V^{*}$ by proving that $L_{Q^{*}}$ is an upper bound on $L_{V^{*}}$:
Proposition 3.
Given a Lipschitz RMDP as in Propositions 1 and 2, $L_{V^{*}}\leq L_{Q^{*}}$.
Proof.
See Appendix A.4.
4.3 Explicit Lipschitz Value Estimation (ELVEn)
In this section, we introduce Explicit Lipschitz Value Estimation (ELVEn), which combines the Fast Gradient Sign Method (FGSM) [32] and Lipschitz regularization to reduce the sensitivity to $\epsilon$ when estimating the infimum in the robust Bellman operator. GBR is closely related to FGSM since
$V(s^{\prime}_{0})-\epsilon||\nabla_{s}V(s^{\prime}_{0})||_{1}=V(s^{\prime}_{0})+\langle\nabla_{s}V(s^{\prime}_{0}),\hat{s}-s^{\prime}_{0}\rangle,\quad\hat{s}=s^{\prime}_{0}-\epsilon\cdot\text{sign}(\nabla_{s}V(s^{\prime}_{0}))$ (10)
When applying FGSM to the critic, we denote the adversarial state as $Adv(s,\epsilon,V)$, which is generated by attacking a critic $V$ at the input state $s$ given a radius $\epsilon$. Then the adversarial computation of $\hat{s}$ in Eq. 10 is precisely FGSM since $\hat{s}=Adv(s^{\prime}_{0},\epsilon,V)$, meaning that GBR is equivalent to evaluating the first-order Taylor expansion of $V$ around $s^{\prime}_{0}$ at $\hat{s}$. However, GBR does not explicitly compute $\hat{s}$ or reevaluate $V(\hat{s})$, but estimates $||\nabla_{s}V(s^{\prime}_{0})||_{1}$, which may be less accurate before training is complete. We hypothesize that explicitly using $V(\hat{s})$ to update the robust Bellman operator, rather than implicitly evaluating the first-order Taylor expansion around $s^{\prime}_{0}$, will provide a more accurate estimate of $\inf_{\bar{s}\in B_{\epsilon}(s^{\prime}_{0})}V(\bar{s})$. This does incur the cost of one additional forward-pass evaluation of $V$ per update, but has the potential benefit of reducing sensitivity to $\epsilon$ and improving robustness.
The foregoing modification still does not mitigate the issue of a large and ill-conditioned $\nabla_{s}V$. To also mitigate this issue, we perform Lipschitz regularization on the critic network based on Propositions 1-3. However, optimizing the training loss subject to a hard Lipschitz constraint is an infeasible problem. It is not only intractable to determine the precise Lipschitz constants in advance, but also computing the Lipschitz constant of neural networks is considered NP-hard [21, 10]. Instead, we simplify the constrained problem by adding a penalty term to the objective that is proportional to the square of
$\frac{|V(s)-V(\hat{s})|}{d_{\mathcal{S}}(s,\hat{s})}=\frac{|V(s)-V(\hat{s})|}{\epsilon}.$ (11)
4.4 ELVEn-SAC
We apply ELVEn as proposed above to Soft Actor-Critic (SAC) [33], a popular off-policy actor-critic algorithm, and call the resulting implementation “ELVEn-SAC.” The implementation is described below and with pseudocode in Algorithm 1. For the policy network $\pi_{\psi}(a|s)$, we use the reparameterization trick when predicting the action distribution. For the critic network, we use two separate networks $Q_{\theta_{i}}(s,a),i=\{1,2\}$ to perform clipped double Q-learning [34] and use two target networks $Q_{\hat{\theta}_{i}}(s,a)$ to soft-update the corresponding critic. In this case we replace $V$ in Sec. 4.3 by $Q$ such that the adversarial state of $Q$ at state $s$ is $Adv(s,\epsilon,Q)=s-\epsilon\cdot\text{sign}(\nabla_{s}Q(s,a))$. The SAC temperature $\alpha$ in ELVEn-SAC is a trainable parameter responsible for adjusting the weight of the entropy term. We denote the input transition data-point as $(s,a,s^{\prime},r)$ and the perturbation radius as $\epsilon$.
The critic loss of ELVEn-SAC consists of two terms: a prediction loss and a Lipschitz regularization loss. Recall that unlike SAC, the worst-case value functions are used to update the robust Bellman operator. Therefore, ELVEn-SAC always calculates the worst-case $Q$-values with the critics and their targets. The learning target $\hat{Q}$ is defined as
$\hat{Q}(s,a)=r+\gamma\mathbb{E}_{a^{\prime}\sim\pi_{\psi}(\cdot|s^{\prime})}\left[\min_{i=1,2}Q_{\hat{\theta}_{i}}(Adv(s^{\prime},\epsilon,Q_{\hat{\theta}_{i}}),a^{\prime})-\alpha\log\pi_{\psi}(a^{\prime}|s^{\prime})\right].$ (12)
The prediction loss minimizes the residual in the robust Bellman operator by learning $\hat{Q}(s,a)$
$\mathcal{L}^{\text{Pred}}_{Q}(\theta)=\left(Q_{\theta}(s,a)-\hat{Q}(s,a)\right)^{2}$ (13)
The Lipschitz regularization loss minimizes the critic network change at adversarial states
$\mathcal{L}^{\text{Lips}}_{Q}(\theta)=\lambda\left(Q_{\theta}(s,a)-Q_{\theta}(Adv(s,\epsilon,Q_{\theta}),a)\right)^{2}$ (14)
where $\lambda$ is a hyperparameter controlling the importance of regularization versus temporal difference learning, and wraps up the Lipschitz regularization coefficient $\frac{1}{\epsilon}$. The final critic loss is defined as
$\mathcal{L}_{Q}(\theta)=\mathcal{L}^{\text{Pred}}_{Q}(\theta)+\mathcal{L}^{\text{Lips}}_{Q}(\theta)$ (15)
The actor loss maximizes the entropy of the policy, relative to the current action value estimates
$\mathcal{L}_{\pi}(\psi)=\mathbb{E}_{\bar{a}\sim\pi_{\psi}(\cdot|s)}\left[\alpha\log\pi_{\psi}(\bar{a}|s)-\min_{i=1,2}Q_{\theta_{i}}(Adv(s,\epsilon,Q_{\theta_{i}}),\bar{a})\right]$ (16)
We tune the temperature parameter $\alpha$ with the following loss, where $\mathcal{H}$ is a constant target entropy:
$\mathcal{L}_{\alpha}(\alpha)=-\alpha\cdot\mathbb{E}_{\bar{a}\sim\pi_{\psi}(\cdot|s)}\left[\log\pi_{\psi}(\bar{a}|s)+\mathcal{H}\right]$ (17)
Algorithm 1 Explicit Lipschitz Value Estimation SAC (ELVEn-SAC)
Critics $Q_{\theta_{1}}(s,a)$, $Q_{\theta_{2}}(s,a)$ and targets $Q_{\hat{\theta}_{1}}(s,a)$, $Q_{\hat{\theta}_{2}}(s,a)$. Policy $\pi_{\psi}(a|s)$. Temperature parameter $\alpha$. Learning rates $\eta_{\theta}$, $\eta_{\psi}$, $\eta_{\alpha}$. Soft-update coefficient $\tau$. Replay buffer $\mathcal{D}$.
Update $Q_{\theta_{1}}(s,a)$, $Q_{\theta_{2}}(s,a)$ and $\pi_{\psi}(a|s)$.
for each time step $t$ do
     $a_{t}\sim\pi_{\psi}(\cdot|s_{t})$, $s_{t+1}=T(s_{t},a_{t})$, $r_{t}=R(s_{t},a_{t})$
     $\mathcal{D}\leftarrow\mathcal{D}\cup(s_{t},a_{t},s_{t+1},r_{t})$
     for each training step do
         $(s,a,s^{\prime},r)\sim\mathcal{D}$
         $\theta_{i}\leftarrow\theta_{i}-\eta_{\theta}\nabla_{\theta_{i}}\mathcal{L}_{Q}(\theta_{i})$, $i\in\{1,2\}$
         $\psi\leftarrow\psi-\eta_{\psi}\nabla_{\psi}\mathcal{L}_{\pi}(\psi)$
         $\alpha\leftarrow\alpha-\eta_{\alpha}\nabla_{\alpha}\mathcal{L}_{\alpha}(\alpha)$
         $\hat{\theta}_{i}\leftarrow(1-\tau)\hat{\theta}_{i}+\tau\theta_{i}$, $i\in\{1,2\}$
     end for
end for
5 Experiments
5.1 Experimental Setup
Parameter Value Parameter Value
Optimizer Adam Reward scale 1.0
Optimizer parameters $\beta_{1}=0.9,\beta_{2}=0.999$ Target entropy $-\dim(\mathcal{A})$
Hidden layer size [256, 256] Exploration steps $1\times 10^{4}$
Learning rate of $\pi,Q,\alpha$ $3\times 10^{-4}$ Mini-batch size 256
Discount factor $\gamma$ 0.99 Replay buffer capacity $1\times 10^{6}$
Soft-update coefficient $\tau$ 0.005 Update frequency 1 update per step
Table 1: Shared hyperparameters for SAC, SC-SAC and ELVEn-SAC. Left: Hyperparameters for MLP actor and critic network. Right: Hyperparameters for policy training.
We conduct experiments on four tasks: Ant, HalfCheetah, Hopper and Walker2d on MuJoCo (version 2.2.0) [35]. We firstly discuss our Lipschitz assumption on these environments. For the reward formulation, all four environments have a consistent form $R(s,a)=\alpha_{1}v+\alpha_{2}-\alpha_{3}\sum_{i}a^{2}_{i}$, where $v$ is the forward velocity, $a_{i}$ is the $i$-th element of the action vector $a$, and $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{3}$ are task-specific non-negative constants. This reward function is Lipschitz thanks to two facts: (1) $v$ is included in the state vector $s$, and (2) the range of actions are $[-1,1]$, meaning that the gradient norm of the smooth function $\sum_{i}a^{2}_{i}$ is upper bounded. For the transition dynamics, these environments are not strictly Lipschitz due to discontinuous velocity changes when the agents make or break contact with the floor. However, the Lipschitz condition can be locally satisfied in the state-action transitions where dynamics are smooth in between isolated contact events. Hence we believe our Lipschitz regularization objective is still reasonable when averaged over transitions in the replay buffer.
For comparison we use two baseline algorithms: SAC [33] and SC-SAC [7]  [^1] https://github.com/MIRALab-USTC/RL-SCPO. For each task and algorithm, we train a policy on the nominal environment $T_{0}$ for one million (1M) steps, and run 4 independent training repetitions with different random seeds. For evaluation, following [7], we created an uncertainty set of $11\times 11=121$ perturbed environments for each task by rescaling the mass and friction of all rigid bodies. We evaluate performance on each perturbed environment at multiple time-points near the end of training: 970K, 980K, 990K and 1M steps. Each performance evaluation uses 10 independent episodes and calculates the net reward attained in each episode. This is a total of $4$ training repetitions per task, each evaluated in $121$ perturbed environments with $4\times 10=40$ evaluations per environment per repetition. Table 1 lists the hyperparameters fixed in our experiments, where “Exploration” means sampling with random actions at the early stage of training. Computational training costs are given in Appendix B.
5.2 Empirical Performance and Robustness
Figure 1: Robustness visualization with $\epsilon=0.005$ for SC-SAC and ELVEn-SAC and $\lambda=0.1$ for ELVEn-SAC. Each $11\times 1$ 11 panel has one grid point per perturbed environment (central pixel is the default environment). Pixel value indicates average net reward for the corresponding perturbed environment across its 160 evaluation episodes; darker red represents higher reward. Pixel-wise differences between paired algorithms are shown in the right half of the figure. In particular, blue in the two rightmost columns indicates where ELVEn-SAC outperformed the baselines.
Figure 1 shows a qualitative comparison of the three algorithms. This figure visualizes performance on each perturbed environment, averaged over the $4$ training repetitions and $40$ evaluation episodes ($160$ episodes total). The figures show that ELVEn-SAC performs similarly to the other algorithms near the nominal environment $T_{0}$, and often favorably in perturbed environments farther from $T_{0}$.
We also computed several metrics on the evaluation data to quantitatively compare the three algorithms, shown in Fig. 2, 3 and Table 2. The first metric is net reward on the default environment $T_{0}$, averaged over the $40$ evaluation episodes near the end of training. Improved robustness usually comes at the cost of reduced performance on the default environment, but we would like this reduction to be small. As Fig. 2 and Table 2 show, ELVEn-SAC performs comparably to, and in one case better than, the other two algorithms on the default environment.
The second metric is net reward averaged over all $121$ perturbed environments, not only the default environment. A robust algorithm should perform reasonably well on all environments in the uncertainty set, in which case this metric should be relatively large. The results show that by this metric, ELVEn-SAC outperforms the others on three of four environments.
Our third metric is more nuanced and designed to closely match the notion of “robustness.” Robustness is concerned with the worst-case performance within a given attack radius. With respect to the $11\times 1$ 11 perturbation grids, we consider attack radii ranging from 0 to 5, measured by max-norm distance from the grid center $(5,5)$ which corresponds to the nominal environment $T_{0}$. Formally, at each grid point $(i,j)$, we first calculate the average net reward $\overline{R}_{i,j}$ over the $40$ evaluation episodes for the corresponding perturbed environment $T_{i,j}$. Then, for a given attack radius $\rho$, we take the minimum over all environments within $\rho$ of the default environment, i.e.:
$\rho\text{-robustness}=\underset{(i,j)\in\mathcal{B}_{\rho}(5,5)}{\text{min}}\overline{R}_{i,j}$ (18)
where $\mathcal{B}_{\rho}(5,5)$ is the max-norm ball of radius $\rho$ centered at grid point $(5,5)$. Finally, to obtain an overall measure of robustness, we average this metric over $\rho\in\{0,1,...,5\}$. Overall robustness is shown in Figure 2 and Table 2. Individual values of $\rho$-robustness for each $\rho$ are shown in Figure 3. The results again show that by this metric, ELVEn-SAC outperforms the others in 3 of 4 tasks, with noticeably higher $\rho$-robustness in Walker2d and Hopper for several values of $\rho$.
Reward in $T_{0}$ Reward across $\mathcal{T}_{\epsilon}$ Robustness Environment SAC SC-SAC ELVEn-SAC SAC SC-SAC ELVEn-SAC SAC SC-SAC ELVEn-SAC Ant 5445 5245 5439 4739 4626 4751 3308 3485 3602 HalfCheetah 8961 9462 9019 7062 7297 6907 6619 6809 6041 Walker2d 4321∗∗ 4844 4825 3426 3723 4198 2924 2781 3649 Hopper 2905 2795 3324 1522∗∗∗ 1617 2132 1266∗∗∗ 1426 1943
Table 2: Performance metrics for each environment and algorithm, averaged over training repetitions; higher is better. Bold-face indicates the best algorithm for each metric and environment. Super-scripts , ∗∗, and ∗∗∗ indicate baseline metrics that were outperformed by ELVEn-SAC within $p=0.1$, $0.05$, and $0.01$ significance levels, respectively.
Finally, to gauge the statistical significance of these results, we used 1-sided Welch’s t-tests for samples with unequal variance, the alternative hypotheses being that ELVEn-SAC was higher than another algorithm for a given metric and task. Since evaluations within a given training repetition are not independent, we calculated the metrics on a per-repetition basis and then treated the four repetitions as a size-4 sample for the hypothesis tests. Due to the small sample size, we had limited statistical power and most of ELVEn-SAC’s performance improvements were not significant at a $p=0.05$ significance level. However, we can note that ELVEn-SAC was never significantly worse than any other algorithm, and in some cases, significantly better than SAC. Some comparisons with SC-SAC are also approaching significance with $p\leq 0.1$.
Figure 2: Performance metrics for each algorithm and environment, one panel per metric. Per-repetition metrics are plotted as scatter points; bar heights are averaged over training repetitions.
Figure 3: $\rho$-robustness values for each algorithm, environment, and attack radius $\rho\in\{0,...,5\}$. Per-repetition values are shown as transparent markers; lines are averaged over training repetitions.
Figure 4: The sensitivity of SC-SAC and ELVEn-SAC to $\epsilon$ on Hopper. We compare the performance when $\epsilon\in\{0.001,0.003,0.005,0.007,0.01\}$. In ELVEn-SAC, $\lambda$ is fixed to 0.1.
Figure 5: The sensitivity of ELVEn-SAC to $\lambda$ on Hopper when $\epsilon\in\{0.005,0.007,0.01\}$. For comparison, we attach the performance of SC-SAC with the same $\epsilon$ in the leftmost column. From the second column to the rightmost, $\lambda=0,0.1,0.2,0.5$.
5.3 Hyper-parameter Analyses
We also empirically analyzed the sensitivity of SC-SAC and ELVEn-SAC to $\epsilon$. Fig. 4 shows the evaluation results of policies trained with different $\epsilon$. When $\epsilon$ is within a specific range, both SC-SAC and ELVEn-SAC achieve good generalization in perturbed environments. For larger $\epsilon$, SC-SAC’s performance drops while ELVEn-SAC’s performance is more stable. Therefore, ELVEn-SAC is less sensitive to the choice of $\epsilon$ than SC-SAC, showing robustness against larger perturbations.
Fig. 5 shows how robustness of ELVEn-SAC can improve and eventually exceed that of SC-SAC when large perturbations occur (large $\epsilon$ values). It shows that ELVEn-SAC without Lipschitz regularization ($\lambda$ = 0) occasionally fails to outperform SC-SAC, and confirms the necessity of Lipschitz regularization in ELVEn-SAC ($\lambda$ > 0). Furthermore, Fig. 6 visualizes the smoothness of the trained $Q$ networks. We observe that $Q$ networks trained by ELVEn-SAC are smoother than the one trained by SC-SAC, and become smoother with more aggressive regularization (higher $\lambda$). Fig. 6 shows different gradient slopes for SC-SAC vs. ELVEn-SAC; we hypothesize this is due to the different update rules for the $Q$ networks.
Performance plots for the full range of $\epsilon$ and $\lambda$ across all four tasks are included in Appendix C.
(a) SC-SAC
(b) $\lambda=0$
(c) $\lambda=0.1$
(d) $\lambda=0.5$
Figure 6: The 3D landscape of predicted $Q(s,a)$ on Hopper for SC-SAC (red) and ELVEn-SAC (green), where $\epsilon$ is fixed to 0.005. The variable $s[0]$ and $s[1]$ denote the first two dimensions of state space. We vary $s[0]$ and $s[1]$, and fix the other states and actions to 0. Note the different z-axis ranges for SC-SAC and ELVEn-SAC.
6 Conclusion
In this paper, we tested our hypothesis that Lipschitz regularization of the value function can enhance robustness to environmental perturbations. We introduced Explicit Lipschitz Value Estimation (ELVEn), a combination of the Fast Gradient Sign Method and Lipschitz regularization, and integrated it into an off-policy algorithm ELVEn-SAC. Our experimental results show that ELVEn-SAC is less sensitive to perturbations than prior work. In future work, we aim to further explore ELVEn in conjunction with local Lipschitz regularization, evaluate ELVEn on a wider range of neural architectures and tasks, and test its utility for sim-to-real transfer on real physical robotic hardware. Furthermore, our Welch’s t-test is currently based on insufficient repetitions, so we aim to improve its reliability by conducting more experiments. In the long term, our work can have positive broader impacts on society since robustness is important in safety-critical RL applications. That said, autonomous systems have inherent risks and potential negative effects on society, which must be considered when our methods are deployed on real hardware.
Appendix
A Proofs
A.1 Lemma 1
See 1
Proof.
For Eq. 6, suppose without loss of generality that $\inf_{x}f(x)\geq\inf_{y}g(y)$ (otherwise, use the following reasoning with $f$ and $g$ swapped). Letting $y^{*}$ denote a minimizer of $g$, we have
$|\inf_{x}f(x)-\inf_{y}g(y)|=\inf_{x}f(x)-g(y^{*})\leq f(y^{*})-g(y^{*})=|f(y^{*})-g(y^{*})|\leq\sup_{z}|f(z)-g(z)|.$ (19)
We know that $y^{*}$ exists since the domain is compact, and we can reintroduce absolute value signs because the the leftmost quantity is non-negative. The reasoning for Eq. 7 is similar. ∎
A.2 Proposition 1
See 1
Proof.
Following [30], we first show Lipschitz continuity for each finite horizon, $n$-step optimal value function (denoted $V_{n}$) by induction on $n$. Then we recover the bound on $L_{V^{*}}$ in the limit $n\rightarrow\infty$. In the base case $n=0$, we have $V_{0}(s)=\max_{a}R(s,a)$, and
$|V_{0}(s_{1})-V_{0}(s_{2})|=|\max_{a_{1}}R(s_{1},a_{1})-\max_{a_{2}}R(s_{2},a_{2})|\leq\max_{a}|R(s_{1},a)-R(s_{2},a)|\leq L_{R}d_{\mathcal{S}}(s_{1},s_{2}),$
where the first inequality uses Lemma 1 and the second uses $d_{\mathcal{SA}}((s_{1},a),(s_{2},a))=d_{\mathcal{S}}(s_{1},s_{2})$.
For the inductive case, we note that the $(n+1)$-step value function satisfies
$V_{n+1}(s)=\max_{a}\left(R(s,a)+\gamma\inf_{T\in\mathcal{T}_{\epsilon}}V_{n}(T(s,a))\right).$ (20)
Assuming the inductive hypothesis that $V_{n}$ is $L_{n}$-Lipschitz continuous, we derive $L_{n+1}$ as follows:
$|V_{n+1}(s_{1})-V_{n+1}(s_{2})|$ (21)
$=$ $\left|\max_{a_{1}}\left(R(s_{1},a_{1})+\gamma\inf_{T_{1}}V_{n}(T_{1}(s_{1},a_{1}))\right)-\max_{a_{2}}\left(R(s_{2},a_{2})+\gamma\inf_{T_{2}}V_{n}(T_{2}(s_{2},a_{2}))\right)\right|$ (22)
$\leq$ $\max_{a}\left|\left(R(s_{1},a)+\gamma\inf_{T_{1}}V_{n}(T_{1}(s_{1},a))\right)-\left(R(s_{2},a)+\gamma\inf_{T_{2}}V_{n}(T_{2}(s_{2},a))\right)\right|$ (23)
$\leq$ $\left|R(s_{1},a^{*})-R(s_{2},a^{*})\right|+\gamma\left|\inf_{T_{1}}V_{n}(T_{1}(s_{1},a^{*}))-\inf_{T_{2}}V_{n}(T_{2}(s_{2},a^{*}))\right|$ (24)
$\leq$ $L_{R}d_{\mathcal{S}}(s_{1},s_{2})+\gamma\sup_{T}\left|V_{n}(T(s_{1},a^{*}))-V_{n}(T(s_{2},a^{*}))\right|$ (25)
$=$ $L_{R}d_{\mathcal{S}}(s_{1},s_{2})+\gamma\left|V_{n}(T^{*}(s_{1},a^{*}))-V_{n}(T^{*}(s_{2},a^{*}))\right|\leq(L_{R}+\gamma L_{n}L_{T^{*}})d_{\mathcal{S}}(s_{1},s_{2}),$ (26)
where $a^{*}$ and $T^{*}$ are the respective maximizers. We used Lemma 1 in Eq. 23 and Eq. 25 and simplified $d_{\mathcal{SA}}$ similarly to the base case. Expanding the recurrence $L_{n+1}\leq L_{R}+\gamma L_{n}L_{T^{*}}$, taking the limit $n\rightarrow\infty$ and using properties of geometric series, the claim follows. ∎
A.3 Proposition 2
See 2
Proof.
$Q^{*}$ and $V^{*}$ are related by
$Q^{*}(s,a)=R(s,a)+\gamma\inf_{T\in\mathcal{T}_{\epsilon}}V^{*}(T(s,a)).$ (27)
Therefore we have
$|Q^{*}(s_{1},a_{1})-Q^{*}(s_{2},a_{2})|$ $\leq|R(s_{1},a_{1})-R(s_{2},a_{2})|+\gamma\sup_{T}|V^{*}(T(s_{1},a_{1}))-V^{*}(T(s_{2},a_{2}))|$ (28)
$\leq(L_{R}+\gamma L_{V^{*}}L_{T^{*}})\cdot d_{\mathcal{SA}}((s_{1},a_{1}),(s_{2},a_{2}))$ (29)
$\leq(L_{R}+\gamma\frac{L_{R}}{1-\gamma L_{T^{*}}}L_{T^{*}})\cdot d_{\mathcal{SA}}((s_{1},a_{1}),(s_{2},a_{2}))$ (30)
$=\frac{L_{R}}{1-\gamma L_{T^{*}}}\cdot d_{\mathcal{SA}}((s_{1},a_{1}),(s_{2},a_{2})),$ (31)
where Eq. 28 uses Lemma 1 yet again and Eq. 30 uses the bound on $L_{V^{*}}$ from Proposition 1. Therefore $L_{Q^{*}}\leq L_{R}/(1-\gamma L_{T^{*}})$. ∎
A.4 Proposition 3
See 3
Proof.
$V^{*}$ and $Q^{*}$ are also related by $V^{*}(s)=\max_{a}Q^{*}(s,a)$. Therefore we have
$|V^{*}(s_{1})-V^{*}(s_{2})|$ $=|\max_{a_{1}}Q^{*}(s_{1},a_{1})-\max_{a_{2}}Q^{*}(s_{2},a_{2})|$ (32)
$\leq\max_{a}|Q^{*}(s_{1},a)-Q^{*}(s_{2},a)|$ (33)
$=|Q^{*}(s_{1},a^{*})-Q^{*}(s_{2},a^{*})|$ (34)
$\leq L_{Q^{*}}d_{\mathcal{S}}(s_{1},s_{2}),$ (35)
where $a^{*}$ is again the relevant maximizer. Once more we have used Lemma 1 in Eq. 33, and simplified $d_{\mathcal{SA}}$ in Eq. 35 similarly to the base case of Proposition 1. Since $s_{1}$ and $s_{2}$ are arbitrary, $|V^{*}(s_{1})-V^{*}(s_{2})|/d_{\mathcal{S}}(s_{1},s_{2})\leq L_{Q^{*}}$ for all $s_{1}\neq s_{2}$. Therefore the Lipschitz constant $L_{V^{*}}\leq L_{Q^{*}}$ also. ∎
B Comparisons of Computational Cost
We follow [7] to implement observation normalization in SAC, SC-SAC and ELVEn-SAC. For each task, we firstly run SAC without observation normalization for 1M steps and compute the mean and standard deviation with the observations in the replay buffer. Then we use these fixed values to normalize the observations. We run all experiments on a shared GPU cluster where the primary GPU model is Quadro RTX 6000, and compare the computation costs in Table 3. However, due to the resource allocation on the shared cluster, making an objective comparison can be challenging.
Task Ant HalfCheetah Hopper Walker2d
SAC 10.096 10.132 9.931 10.813
SC-SAC 17.916 21.429 20.018 19.629
ELVEn-SAC 16.195 16.035 15.697 15.797
Table 3: The estimated time (unit: ms) of updating 1 batch (256 samples).
C Complete Experimental Plots
We report our experimental results of baselines (SAC and SC-SAC) as follow.
The experimental results of ELVEn-SAC are attached. For each hyperparameter setting ($\epsilon$ and $\lambda$), we compare the corresponding ELVEn-SAC to SAC, and SC-SAC with the same $\epsilon$ by following the strategy in Fig. 1.